# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v1

@@ I Introduction and main results
Efficient sources of single and indistinguishable photons Eisaman et al. (2011); Senellart et al. (2017); Slussarenko and Pryde (2019) are a fundamental requirement to perform all kinds of quantum information tasks with photons: photonic quantum computation Knill et al. (2001); Kok et al. (2007); Tiecke et al. (2014); Tiarks et al. (2016); Takeda and Furusawa (2019), networking Lodahl (2017), simulation Aspuru-Guzik and Walther (2012); Hartmann (2016), communication Briegel et al. (1998); Sangouard et al. (2011); Llewellyn et al. (2020), cryptography Gisin et al. (2002), or boson sampling Brod et al. (2019); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a). Scaling up these protocols requires the generation of a large product-state of $N\gg 1$ indistinguishable photons,
@@
The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al. (2016), which can be deterministically excited He and Barkai (2006a, b) and decay spontaneously, producing individual photons that are collected into the desired channels. The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs. A long list of setups includes single atoms McKeever et al. (2004); Hijlkema et al. (2007), single molecules Brunel et al. (1999); Ahtee et al. (2009); Rezai et al. (2018), trapped ions Almendros et al. (2009); Stute et al. (2012); Higginbottom et al. (2016); Sosnova et al. (2018), and atomic ensembles Chou et al. (2004); Matsukevich et al. (2006); Farrera et al. (2016); Ripka et al. (2018), as well as solid-state systems Aharonovich et al. (2016) such as quantum dots Santori et al. (2002); Nawrath et al. (2019); Wang et al. (2019b); Schnauber et al. (2019); Thoma et al. (2016); Dusanowski et al. (2019); Kirsanske et al. (2017); Liu et al. (2018); Uppu et al. (2020) or color centers in diamond Rodiek et al. (2017); Sipahigil et al. (2014); Alléaume et al. (2004); Gaebel et al. (2004); Wu et al. (2007); Babinec et al. (2010); Tchernij et al. (2017); Neu et al. (2011); Schröder et al. (2011); Wang et al. (2019c). In the microwave regime, superconducting quantum circuits have been used to create nearly deterministic SPSs Houck et al. (2007); Pechal et al. (2014); Peng et al. (2016); Forn-Díaz et al. (2017); Gasparinetti et al. (2017); Pfaff et al. (2017); Eder et al. (2018); Zhou et al. (2020), which have the advantage of being externally tuneable Peng et al. (2016); Forn-Díaz et al. (2017); Pfaff et al. (2017), fast Pechal et al. (2014); Zhou et al. (2020), and readily integrated on-chip with very low losses Zhou et al. (2020). Currently, quantum dots in micro-pillar cavities are proving the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al. (2019c); Uppu et al. (2020), but it is hard to manufacture many of them identically, and tuning them also remains elusive for more than two emitters Patel et al. (2010); Flagg et al. (2010); Moczała-Dusanowska et al. (2019); Ellis et al. (2018); Kambs and Becher (2018).
@@ II.1 General cavity-QED model
Let us consider $N$ two-level systems or “qubits” coupled to an optical or microwave cavity as shown in Figs. 1(a) and (b). The cavity mode of frequency $\omega_{c}$ is coherently driven by a time-dependent field with frequency $\omega_{d}$, amplitude $\Omega(t)$ and detuning $\Delta=\omega_{d}-{\color[rgb]{0,0,0}\omega_{c}}.$ In a frame rotating with frequency $\omega_{d}$, the Hamiltonian of this system reads
$H(t)={}$ $-\Delta a^{{\dagger}}a+\Omega(t)(a+a^{{\dagger}})$
$-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}a).$ (4)
Here, $^{{\dagger}}, aa$ are the cavity creation and annihilation operators, whereas $\sigma_{j}^{+}=|e\rangle_{j}\langle g|$, $\sigma_{j}^{-}=|g\rangle_{j}\langle e|$, and $\sigma_{j}^{z}=|e\rangle_{j}\langle e|-|g\rangle_{j}\langle g|$ are standard Pauli operators for qubits $j=1,\dots,N$, with ground and excited states denoted by $|g\rangle_{j}$, and $|e\rangle_{j}$, respectively. Each qubit has a possibly different frequency $\omega_{q}^{j}$ and detuning $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ with respect to the cavity drive. The interaction strength $g_{j}$ between the $j$-th qubit and the cavity satisfies the rotating-wave approximation (RWA) $g_{j}\ll\omega_{q}^{j}\sim\omega_{d},$ and coupling terms can be written in the standard Jaynes-Cummings form. We control the emitters by modulating the envelope of the cavity drive
@@ II.2 Cavity-mediated synchronization and residual correlations
The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode, using a strong drive $|\Omega_{0}|\gg\kappa,|\Delta|$. We model this state using a displacement of the Fock operator
@@
$\dot{\alpha}(t)={}$ $-\left[\kappa/2-i\Delta\right]\alpha(t)-i\Omega_{0}f(t).$ (8)
Once the pulse is switched on ($f(t)=1$), $|\alpha(t)|$ grows and reaches the steady state value
@@
As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, described by the effective Hamiltonian, $\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j}g_{j}|\alpha(t)|(\sigma_{j}^{-}+\sigma_{j}^{+}).$ In the asymptotic limit, the large and steady cavity amplitude allows for a strong cavity-mediated driving strength on the qubits
$\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (10)
which we use to excite all SPSs simultaneously.
In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed. To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
$g_{j},|\delta_{j}|\ll\kappa,|\Delta|.$ (11)
These conditions [cf. App. A.2] ensure that the cavity reaches a steady state (9) where quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated. The dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\},$ is then modeled by the master equation
@@
with the effective Hamiltonian $\tilde{H}(t)$
@@
In addition to the cavity-mediated driving $\Omega_{\rm cm}^{j}$ in Eq. (10), the qubits experience a Lamb-shift
@@ II.3 Synchronization dynamics and parameter regime
Third, to achieve nearly independent photons in a product state (1), the photons must be created faster than the speed of the cavity mediated correlations and interactions ($\gamma\gg J_{\rm cm},\gamma_{\rm cm}$). This is the only intrinsic limitation of the proposed multiphoton scheme. In Sec. III-IV we show that state-of-the-art circuit-QED setups satisfy this requirement and allow the synchronization of $N\sim 100-1000$ nearly independent SPSs.
@@
Figure 2: Synchronization of two deterministic SPSs via a cavity. (a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity [cf. red dash-dotted line]. This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf. blue solid line]. After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$. (b) Long time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its own antenna [cf.  green dashed line]. Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. orange dotted line]. The parameters correspond to a state-of-the-art circuit-QED implementation, and are presented in row C of Table 1.
@@
Fig. 2 illustrates the two processes in the cavity-mediated synchronization method, using two sychronized SPSs with parameters that satisfy Eq. (17). Fig. 2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|\sim 10^{4}$ [cf. red dash-dotted line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf. blue solid line]. Fig. 2(b) shows the long time dynamics, in which the photons are generated. Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf. blue solid line]. To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs. III.2-III.3. For non-optimized parameters satisfying Eq. (17), we reach ${\cal P}_{1}\approx 0.99$ [cf. green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. orange dotted line]. This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq. (2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$). We remark that this example only illustrates the working principle of the setup. As we show in Secs. III and IV, the parameters can be further optimized to improve the efficiency and scalability of the multi-photon emitter.
@@ III.1 State-of-the-art parameters
In Fig. 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits. For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency on the order of a few GHz. These superconducting resonators support strong drives and can hold a large number of photons in steady state Pietikäinen et al. (2017); Jenkins et al. (2014); Macri et al. (2020). In our design we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al. (2017); Jenkins et al. (2014).

### v2

@@ I Introduction and main results
Efficient sources of single and indistinguishable photons Eisaman et al. (2011); Senellart et al. (2017); Slussarenko and Pryde (2019) are a fundamental requirement to perform all kinds of quantum information tasks with photons: photonic quantum computation Knill et al. (2001); Kok et al. (2007); Tiecke et al. (2014); Tiarks et al. (2016); Takeda and Furusawa (2019), networking Lodahl (2017), simulation Aspuru-Guzik and Walther (2012); Hartmann (2016), communication Briegel et al. (1998); Sangouard et al. (2011); Llewellyn et al. (2019), cryptography Gisin et al. (2002), metrology Motes et al. (2015); Ge et al. (2018), boson sampling Brod et al. (2019); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a), or even quantum optical neural networks Steinbrecher et al. (2019). Scaling up these protocols requires the generation of a large product-state of $N\gg 1$ indistinguishable photons,
@@
The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al. (2016), which can be deterministically excited He and Barkai (2006); He (2006) and decay spontaneously, producing individual photons that are collected into the desired channels. The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs. A long list of setups includes single atoms McKeever et al. (2004); Hijlkema et al. (2007), single molecules Brunel et al. (1999); Ahtee et al. (2009); Rezai et al. (2018), trapped ions Almendros et al. (2009); Stute et al. (2012); Higginbottom et al. (2016); Sosnova et al. (2018), and atomic ensembles Chou et al. (2004); Matsukevich et al. (2006); Farrera et al. (2016); Ripka et al. (2018), as well as solid-state systems Aharonovich et al. (2016) such as quantum dots Santori et al. (2002); Nawrath et al. (2019); Wang et al. (2019b); Schnauber et al. (2019); Thoma et al. (2016); Dusanowski et al. (2019); Kiršanskė et al. (2017); Liu et al. (2018); Uppu et al. (2020) or color centers in diamond Rodiek et al. (2017); Sipahigil et al. (2014); Alléaume et al. (2004); Gaebel et al. (2004); Wu et al. (2007); Babinec et al. (2010); Tchernij et al. (2017); Neu et al. (2011); Schröder et al. (2011); Wang et al. (2019c). In the microwave regime, superconducting quantum circuits have been used to create nearly deterministic SPSs Houck et al. (2007); Pechal et al. (2014); Peng et al. (2016); Forn-Díaz et al. (2017); Gasparinetti et al. (2017); Pfaff et al. (2017); Eder et al. (2018); Zhou et al. (2020), which have the advantage of being externally tuneable Peng et al. (2016); Forn-Díaz et al. (2017); Pfaff et al. (2017), fast Pechal et al. (2014); Zhou et al. (2020), and readily integrated on-chip with very low losses Zhou et al. (2020). Currently, quantum dots in micro-pillar cavities are proving the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al. (2019c); Uppu et al. (2020), but it is hard to manufacture many of them identically, and tuning them also remains elusive for more than two emitters Patel et al. (2010); Flagg et al. (2010); Moczała-Dusanowska et al. (2019); Ellis et al. (2018); Kambs and Becher (2018).
@@ II.1 General cavity-QED model
Let us consider $N$ two-level systems or “qubits” coupled to an optical or microwave cavity as shown in Figs. 1(a) and (b). This cavity mode of frequency $\omega_{c}$ is externally driven by a coherent field with time-dependent amplitude $\Omega(t)$ and frequency $\omega_{d}$. The total Hamiltonian of the system reads
$H(t)={}$ $\omega_{c}a^{{\dagger}}a+2\Omega(t)\cos(\omega_{d}t)(a+a^{{\dagger}})$
$+\frac{1}{2}\sum_{j=1}^{N}\omega_{j}^{q}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(a+a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+}).$ (4)
Here, $^{{\dagger}}, aa$ are the cavity creation and annihilation operators, whereas $\sigma_{j}^{+}=|e\rangle_{j}\langle g|$, $\sigma_{j}^{-}=|g\rangle_{j}\langle e|$, and $\sigma_{j}^{z}=|e\rangle_{j}\langle e|-|g\rangle_{j}\langle g|$ are standard Pauli operators for qubits $j=1,\dots,N$, with ground and excited states denoted by $|g\rangle_{j}$, and $|e\rangle_{j}$, respectively. Each qubit $j$ has a possibly different frequency $\omega_{q}^{j}$ and a coupling $g_{j}$ to the common cavity mode. Note that we do not assume rotating wave approximation (RWA) to allow for a large cavity occupation $\langle a^{\dagger}a\rangle\gg 1$. We control the emitters by modulating the envelope of the cavity drive
@@ II.2 Cavity-mediated synchronization and residual correlations
The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode. This can be achieved via a very strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning. We model the resulting state using a displacement of the Fock operator
@@
$\dot{\alpha}(t)={}$ $-\left[\kappa/2+i\omega_{c}\right]\alpha(t)-2i\Omega_{0}f(t){\color[rgb]{0,0,0}\cos(\omega_{d}t)}.$ (8)
Once the pulse is switched on ($f(t)=1$), $|\alpha(t)|$ grows and stabilizes around the steady state value
@@
As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
[…]
This is the mechanism that allows us to excite all SPSs simultaneously. In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed. To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
$g_{j},|\delta_{j}|\ll\kappa,|\Delta|{\color[rgb]{0,0,0}\ll\omega_{j}^{q},\omega_{c}},$ (11)
where $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ is the detuning between qubit $j$ and cavity drive. These conditions ensure that the cavity reaches a steady state (9), where small quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated and fast oscillations neglected. As shown in App. A.2, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
@@
with the effective Hamiltonian $\tilde{H}(t)$ given by
@@
Here, the cavity-mediated driving $\Omega_{\rm cm}^{j}$ reads
$\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (14)
with the steady state coherent amplitude given in Eq. (9). In addition, the qubits experience a Lamb-shift
@@ II.3 Synchronization dynamics and parameter regime
Third, to achieve nearly independent photons in a product state (1), the photons must be created faster than the speed of the cavity mediated correlations and interactions ($\gamma\gg J_{\rm cm},\gamma_{\rm cm}$). This is the only intrinsic limitation of the proposed multiphoton scheme, and its effect can be mitigated in the case of a bad cavity $g\ll\kappa$ with strong coherent drive $|\alpha_{\rm ss}|\gg 1$. In Sec. III-IV we show that state-of-the-art circuit-QED setups satisfy this requirement and allow the synchronization of $N\sim 100-1000$ nearly independent SPSs.
@@
Figure 2: Synchronization of two deterministic SPSs via a cavity. (a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity with fast residual oscillations due to non-RWA terms [cf. red solid and black dashed lines]. This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf. blue dash-dotted line]. After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$. (b) Long time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its own antenna [cf.  green dashed line]. Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. (c)-(d) Deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|$ between the exact (6) and effective (12) models for the qubit occupations $x=\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$, and photon generation probabilities $x={\cal P}_{1},P_{2}$ as a function of time. We observe great agreement and no heating effects despite the large cavity driving strength $\Omega_{0}=4\omega_{d}\gg\kappa,|\Delta|$. These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row C of Table 1 and discussed in Sec. III.1.
@@
Fig. 2(a)-(b) illustrates the two processes in the cavity-mediated synchronization method, using two synchronized SPSs with parameters that satisfy Eq. (18). Fig. 2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ [cf. red solid line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf. blue dash-dotted line]. Fig. 2(b) shows the long time dynamics, in which the photons are generated. Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf. blue dash-dotted line]. To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs. III.2-III.3. For non-optimized parameters satisfying Eq. (18), we reach ${\cal P}_{1}\approx 0.99$ [cf. green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq. (2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
We performed the numerical simulations in Figs. 2(a)-(b) using both the full model in Eqs. (4)-(6) and the effective model in Eqs. (12)-(17). They show an excellent agreement with deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf. Fig. 2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim 10^{4}\gg 1$. This demonstrates that the strong cavity driving $\Omega_{0}\gg\kappa,|\Delta|$ does not introduce heating in the system, but only creates a large coherent state in the cavity with some fast oscillations around the mean value $|\alpha_{\rm ss}|^{2}\gg 1$ [cf. solid red and dashed black lines in Fig. 2(a)]. Since this fast oscillations have no detrimental impact in the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (12) which uses the mean steady value (9) for describing the effect of the cavity field (see App. A for more details on the approximations involved in the model).
@@ III.1 State-of-the-art parameters
In Fig. 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits. For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz. These superconducting resonators support strong drives and can hold a large number of photons in steady state $|\alpha_{\rm ss}|^{2}\gg 1$ Pietikäinen et al. (2017); Jenkins et al. (2014); Macrì et al. (2020), as it is required in our scheme. For this, we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al. (2017); Jenkins et al. (2014).